OpenStudy (anonymous):

What is the restriction on the quotient of quantity 3 x squared plus 12 x plus 9 divided by quantity x squared minus 4 divided by quantity 4 x plus 4 divided by quantity x plus 2? x ≠ -2 x ≠ -1 x ≠ 1 x ≠ 2

OpenStudy (anonymous):

@whpalmer4 @keke_luvs_u @lovatic4life @Partycool @XcoltenX @Spacelimbus @GTMclachlan @Green_Plan

OpenStudy (anonymous):

@poopsiedoodle @Power2Knowledge @HelpBlahBlahBlah @Mr.ClayLordMath @ChelseaTheRULER @Xmoses1 @zepdrix

OpenStudy (anonymous):

@fc3857 @nepurrta @Jonathan1997 @JA1 @secret66

OpenStudy (anonymous):

@hba @heather040200 @hlee13 @ash2326 @austinL @Ashleyisakitty

OpenStudy (whpalmer4):

Please write your equation using the equation editor button rather than making us puzzle out what it looks like...

OpenStudy (anonymous):

ok i will hold on

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

i tried doing this but i kept getting answers that wasnt an answer choice

OpenStudy (whpalmer4):

\[\frac{\frac{3x^2+12x+9}{x^2-4}}{\frac{4x+4}{x+2}} = \frac{3x^2+12x+9}{x^2-4}*\frac{x+2}{4x+4}\] Restrictions on the expression are those values of \(x\) such that the denominator equals 0. Be sure to do this before you simplify, because the restrictions apply even if the terms that supply them are cancelled out during simplification!

OpenStudy (whpalmer4):

What do you get when you factor the numerators and denominators of those fractions?

OpenStudy (whpalmer4):

(only denominators matter for this question, but you can probably use the practice :-)

OpenStudy (anonymous):

wouldnt the denominator be 4(x+1)

OpenStudy (whpalmer4):

Well, that's the denominator for the fraction on the right. How about the one on the left?

OpenStudy (anonymous):

ohh ummm....

OpenStudy (anonymous):

i cant really see the fraction on the left

OpenStudy (whpalmer4):

You can't see how to factor \(x^2-4\)?

OpenStudy (whpalmer4):

Here's what we have after doing the division: \[\frac{3x^2+12x+9}{x^2-4}*\frac{x+2}{4x+4}\]

OpenStudy (whpalmer4):

(we converted the division into a multiplication)

OpenStudy (anonymous):

(x-2)(x+2)

OpenStudy (whpalmer4):

Very good! So the combined denominator of our multiplication will be \[4(x+1)(x-2)(x+2)\]What are the restrictions on the variable?

OpenStudy (mathmale):

"What are the restrictions on the variable?" And from where do they stem?

OpenStudy (anonymous):

2 right

OpenStudy (mathmale):

Mind explaining where that came from? Is that the only restriction on x?

OpenStudy (mathmale):

"wouldnt the denominator be 4(x+1)?" Yes, PGR! Given that this is in the denom., what can we conclude about restrictions on x?

OpenStudy (whpalmer4):

Don't forget what I said earlier: "Restrictions on the expression are those values of x such that the denominator equals 0."

OpenStudy (anonymous):

oh i really dont know

OpenStudy (anonymous):

but was i right its 2

OpenStudy (whpalmer4):

\[4(x+1)(x-2)(x+2)=0\]Solve for \(x\)

OpenStudy (whpalmer4):

You're only 1/3 right :-)

OpenStudy (anonymous):

darn

OpenStudy (whpalmer4):

For a product of terms to equal 0, one or more of the terms must equal 0, right? So we need to solve \[x+1=0\]\[x-2=0\]\[x+2=0\]3 different values of \(x\), all will make the denominator 0.

OpenStudy (anonymous):

x=-1 x=2 x=-2

OpenStudy (mathmale):

I like whpalmer's approach. He's urging you to determine every one of the factors of the denominator (there are three here), and then to set each factor = to 0, to find the restrictions on the input variable, x. (There are more complicated situations than this, but concentrate on those 3 factors in the denom. for now.)

OpenStudy (anonymous):

but im confused what is the answer though

OpenStudy (whpalmer4):

The answer is all 3 values of \(x\) that make the denominator = 0. What are they?

OpenStudy (mathmale):

Please go back and re-read what whp has advised you to do along the way to finding those answers. Not answer, but answers (there are three). I hope you're taking notes on this discussion, as you'll need to understand and use these tactics in the future when dealing with similar problems.

OpenStudy (whpalmer4):

If those values were \(x=1,x=2.x=3\), it would be written\[x\ne1,x\ne2,x\ne3\] because \(x\) is not allowed to be any of those values.

OpenStudy (anonymous):

but thats not an answer choice

OpenStudy (whpalmer4):

what I wrote was an example of how one notates a function with restrictions on the variable: \[y = f(x), \,x\ne1, x\ne 2, x\ne3\]

OpenStudy (whpalmer4):

I said "if those values were" which was the tip-off that I was not giving you the answer to your problem.

OpenStudy (whpalmer4):

But why don't you tell us what the answer choices you have are...

OpenStudy (mathmale):

There are various ways to write the SET of answers. Whereas whpalmer has written x=-1, x=2, x=-2 (which is correct), one could also write (in set notation) {-1,2,-2}. Try to find a connection between these results and what you and whp have arrived at here.

OpenStudy (anonymous):

i already did

OpenStudy (mathmale):

And what did you find?

OpenStudy (anonymous):

but im learning how how to do it but i want you to make sure that i am doing it right ok

OpenStudy (mathmale):

I understand that, but at some point we all have to take responsibility for finding and checking our answers unaided. So, once again, please share with us the four possible choices of answers.

OpenStudy (anonymous):

look its at the top where you read the question...i had posted it already

OpenStudy (mathmale):

Very good! So the combined denominator of our multiplication will be 4(x+1)(x−2)(x+2) What are the restrictions on the variable? Is it conceivable that you could mark / check off more than 1 of the given answer choices? As before, I believe the correct answer choices are {-1, 2, -2}. Why don't you try to check all three of those?

OpenStudy (mathmale):

And leave x=1 unchecked.

OpenStudy (anonymous):

thats what i am doing right now im replacing X with those three numbers

OpenStudy (mathmale):

OK. If you still feel stuck, I'd suggest you e-mail, call or visit your instructor, showing him/her what you have done so far.

OpenStudy (anonymous):

thats what i am suppose to do right

OpenStudy (anonymous):

@whpalmer4 im doing it right by replacing x with those three numbers

OpenStudy (mathmale):

I strongly urge you to set "becoming less dependent on others for answers" as a goal. I really don't mean to be unkind, but your constant asking for confirmation does not strike me as healthy, and it makes me very uncomfortable. I've suggested that you talk with your teacher in cases like this. Neither whpalmer nor I grade your work; your teacher does. So, would you please consider seeing your teacher if you're not satisfied with the responses/confirmation you're getting or not getting from whpalmer and me.

OpenStudy (mathmale):

I honestly believe that whp and I have done all we can to help you with this particular problem, and that we have given it (and you) our best efforts. Please: move on.

OpenStudy (anonymous):

i been moved what are you talking about i have already found the answer your still sitting on here writing your post here as you seen i stop looking at

OpenStudy (anonymous):

this particular question and moved on working on my other question

OpenStudy (mathmale):

You could have told me that earlier. Your point is what, exactly?

OpenStudy (anonymous):

well i thought i did but now since you know im gone

OpenStudy (anonymous):

what wsa the answer?

OpenStudy (anonymous):

was

OpenStudy (kaylaprincess):

This was obviously her first time doing this, she deserved the confirmation to know if she was doing her work correctly. She seemed to actually be doing the work and learning, as some people here, do not.

OpenStudy (kaylaprincess):

\[x \neq 1 \]